OpenStudy (anonymous):

HELP ME PLEASE!!! In 1990 the population of Alabama was 4,040,587. In 200, the population was 4,447,100. Find the percent of change from 1990 to 2000.

OpenStudy (anonymous):

Percent change = (Difference in Population / Original Population) x 100 Here we know that the difference in population is 4,447,100 - 4,040,587 = 406, 513 The original population, the one we started off with, was 4,040,587. So it follows that: Percent change = (406,513 / 4,040,587) x 100 = 10.1 % Therefore, the population of Alabama increased by 10.1 % since 1990 till 2000.

OpenStudy (anonymous):

@foreverkamixo